OpenStudy (anonymous):

A cube has mass 60g and lenght of edge 2cm. Calculate the inertial angular momentum of the cube along an axis that passes through one of the cube edges.Need all solution

OpenStudy (anonymous):

You can consider the cube as made of many flat squares. So, find the moment of inertia about an axis passing through a end of our imaginary square. Getting the idea?

OpenStudy (anonymous):

nope, can you solve it @koushik_ksv

OpenStudy (anonymous):

Fine, unless you get the idea my working may seem wrong. First about an axis passing through center, I = \[I = \frac{ ml^2 }{6 }\] Then from parallel axes theorem, about and axis passing through end \[I = \frac{ ml^2 }{ 6 } + m (\frac{ l }{\sqrt{2} })^2\] Now,likewise for all squares and adding put m=60/1000 kg l=2/100 m I am getting 0.000016 kg-m^2.